# Comparing Human-Centered Language Modeling: Is it Better to Model Groups, Individual Traits, or Both?

### v1

@@ Abstract
Natural language processing has made progress in incorporating human context into its models, but whether it is more effective to use group-wise attributes (e.g., over-45-year-olds) or model individuals remains open. Group attributes are technically easier but coarse: not all 45-year-olds write the same way. In contrast, modeling individuals captures the complexity of each person’s identity. It allows for a more personalized representation, but we may have to model an infinite number of users and require data that may be impossible to get. We compare modeling human context via group attributes, individual users, and combined approaches. Combining group and individual features significantly benefits user-level regression tasks like age estimation or personality assessment from a user’s documents. Modeling individual users significantly improves the performance of single document-level classification tasks like stance and topic detection. We also find that individual-user modeling does well even without user’s historical data.
Comparing Human-Centered Language Modeling:
Is it Better to Model Groups, Individual Traits, or Both?
[…]
@@ 1 Introduction
Language varies between people. Two strands of human-centered NLP work have modeled the humans behind the language. The first focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group. These attributes include socio-demographics like age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more. The second strand focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
@@
While these two strands have advanced human-centered NLP (Soni et al., 2023), their relative strengths, complementarity, and impact over different tasks are poorly understood. People are not defined by their group membership alone Orlikowski et al. (2023), but individual traits might not generalize. This paper compares the two approaches and their combination in the same framework and tasks in pre-trained large language models. We use the monolingual socio-demographically adapted model from Hung et al. (2023) and the HaRT model from Soni et al. (2022) for the first two types. We use a multi-task learning setup to create GRIT, a PLM based on HaRT, trained with both individual and group human context in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness). We test all systems on five user- and document-level tasks. Note that because we focus on conceptually comparing group and individual traits, we cannot compare to GPT4, which does not support stratification to either attribute.
@@
PLMs trained on individuals and groups enhance user-level regression tasks like age estimation and personality assessment from user’s multiple documents. Such user-level tasks focus on individual person, and our findings show that these people are best modeled as a combination of their groups and individual traits conforming with the notion that a person is a mix of their group attributes and unique characteristics. Document-level categorization tasks, like stance detection, are more personal. A PLM taught within an individual human context alone improves our considered document-level classification tasks.
@@
By their very nature, models of this kind touch upon sensitive user information. For this reason, we take a responsible release strategy, making only the code for the comparisons publicly available and the exact splits of the TrustPilot and Stance datasets used. We build on top of the publicly available code for HaRT and Hung et al. (2023). We acquired the model and data in a secure manner from the authors of Soni et al. (2022) and TrustPilot data splits from the authors of Hung et al. (2023). For more information about the model and data, see Sections 3 and 4. For a discussion of the ethical implications of the model and data, see Section Ethical Considerations.
@@
Our contributions are: (1) We provide an analysis with a comparison of modeling individual users, group socio-demographic features, and both group and individual traits in PLMs. (2) We evaluate the three modeling strategoies on five downstream tasks: two user-level (personality-openness evaluation and age estimation) and three document-level classification tasks (stance detection, topic detection, and age category prediction). (3) We find that user-level regression tasks like estimating age and assessing personality from user’s multiple documents perform better with mixed individual and group human context (GRIT) and document-level tasks like stance detection and topic detection perform better with individual context alone (HaRT).
@@ 2 Integrating Human Context in PLMs
For our comparative study, we use three systems representing the paradigms of human context modeling to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts.
Training with group context.
@@
We build on Hung et al. (2023)’s work to explore demographic adaptation in transformer-based PLMs. They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author. They use the multilingual reviews dataset with demographic labels from Trustpilot [^1] https://www.trustpilot.com/ Hovy (2015). They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection. In our study, we use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4). We use the results from Hung et al. (2023) with the monolingual PLM model BERT on a US-based English dataset with out-of-domain demographic (age) specialization for our comparison study. Out-of-domain data is the Blogs authorship corpus Schler et al. (2006), and in-domain means Trustpilot corpus. To be consistent and fair in comparing with other human context training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
Training with individual human context.
@@
Soni et al. (2022) introduced human language modeling (HuLM), i.e., training a regular language modeling task but including a dynamic individual human context vector derived from the authors’ texts. This vector captures the human states in which the text is generated to induce coherence between different texts generated by the same author. Soni et al. also view this vector as representing the text-derived human factors. They introduce a Human-aware Recurrent Transformer (HaRT), a unidirectional autoregressive PLM that trains for the HuLM task. They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks. We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
Training with both group and individual human context.
We train a PLM that can integrate the author’s individual and group human context knowledge when training for language modeling. We extend Soni et al.’s HaRT model by training for HuLM and predicting a human group attribute in a dynamic muti-task learning setup as used by Hung et al.. We want to induce the individual human context through the author’s language and inject the group context by predicting a group attribute of the author. Predicting group attributes during training can also be viewed as a regularizer for the model, as it constrains the possible output. We discuss the model in detail in section 3. We compare two user-level tasks and a document-level task, as in Soni et al. (2022), and on topic detection and age prediction tasks, as in Hung et al. (2023).
@@ 3 Models
This section describes the models we compare to represent group and individual human contexts.
3.1 Modeling individual human context
@@
where, $w_{t,i}$ is the $i^{th}$ token in the $t^{th}$ block ($B_{t}^{(a)}$) for user $a$. The tokens from the previous blocks are represented using HaRT’s recurrently updated user-state vector.
3.2 Modeling group human context
@@
Hung et al. (2023) train for domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, they adopt a dynamic MTL objective for training with group human context. Kendall et al. interpret homoscedastic uncertainty as task-dependent weighting and derive a multi-task loss function that can optimally learn the weights to balance the impact of multiple loss functions. This approach accounts for the different scales across multiple losses.
@@
3.3 Modeling both individual and group human context
@@
GRIT incorporates both individual and group human contexts using a multi-task learning setup with HaRT that also predicts a socio-demographic attribute of the author in each forward pass using the average of user-state vectors from each non-padded block of the user’s temporally ordered text.
Multi-Task Learning.
GRIT is pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup. The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
@@
We train one model for the attribute age (GRITage) and one for the attribute personality type openness (GRITope), respectively. The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, we adopt a dynamic MTL objective as Hung et al. (2023). We use the joint loss for a continuous and discrete output as derived in Kendall et al. (2018) and compute our joint objective as follows:
@@
3.4 Transfer Learning
To assess the efficiency of GRIT to transfer learning from predicting one group human attribute to another, we experiment with continuing pre-training for each group attribute. To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness). We do the same for GRITope and continued MTL with predicting age.
@@ 4 Experiments
We compare the performance of PLMs adapted to socio-demographic group factors, individual human contexts, and both individual and group contexts. We use the socio-demographic adapted BERT models: BERTage-MLM and BERTDS from Hung et al. (2023), HaRT Soni et al. (2022), and GRIT models, respectively. We use GPT-2HLC from Soni et al. (2022) as a PLM adapted to the social media domain but devoid of human context. We evaluate performances on two user-level regression tasks: predicting age and a personality score (openness), and on three document-level tasks: stance detection, topic detection, and age classification.
@@ 4.1 User Level Regression Tasks
We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment. These tasks require continuous outcomes derived from multiple documents written by a user. We use a subset of the data from consenting Facebook users who shared their demographic and personality scores Kosinski et al. (2012); Park et al. (2015) along with their Facebook posts. This data is essentially the same as GRIT’s pre-training data. However, the test set is from Park et al. (2015) on which HaRT and GPT-2HLC are evaluated.
GRIT is pre-trained on a multi-task learning setup, including predicting a continuous socio-demographic group attribute. GRITage is trained to predict age, and GRITope is trained to predict openness. We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively. We further evaluate these models on the test sets for personality assessment and age estimation after continuing training for these tasks, as described in section 3.4.
@@
Age Estimation
The training and development datasets are identical to the pre-training data for GRIT. Age was self-reported and limited to users under 65. We compare the performance of the models on a test set of 5000 users from Park et al. (2015) and report Pearson correlation ($r$).
Personality Assessment
As for age estimation, the training and development datasets are the same as the pre-training data for GRIT. We compare the performance of the models to predict openness (one’s tendency to be open to new ideas) on a test set of 1943 users and report disattenuated Pearson correlation ($r_{dis}$) metric to account for questionnaire reliability as in Soni et al. (2022).
@@ 4.2 Document-Level Classification Tasks
We use the same subset of the TrustPilot dataset as for topic detection and the same train, development, and test splits from Hung et al. (2023). Given a single review, this task predicts the age group binary label. Age categories are equally represented in each set. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
@@ 5 Results and Discussion
We report results for all the tasks here, discussing their respective impacts on adapting PLMs to individual human context, group context, and both individual and group context.
@@ 5.1 Comparisons Study
User-Level Tasks.
Table 1 shows the two user-level task results. For computational reasons, both HaRT and GRIT use just 4 blocks of training data for these tasks. We find that the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality. In addition, GRIT performs better on the other user-level task using transfer learning and continued training than by fine-tuning it on the averaged user states. Comparing the transfer learning results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find that training PLMs with individual and group human context benefits such multi-document user-level regression tasks. Consequently, we may view the group attribute prediction as a regularizer for the model.
@@
GPT-2HLC
None
0.839
0.521
HaRT
Individual
0.868
0.619
GRITage
Ind + Grp
$\bf{0.890}^{\dagger}$
$\bf{0.658}^{\dagger}$
GRITope
Ind + Grp
$0.884^{\dagger}$
$0.643^{\dagger}$
@@
Document-Level Tasks.
GPT-2HLC
None
68.6
$\bf{69.8}$
65.4
63.9
BERTage-MLM
Group
68.4
64.6
$61.9^{*}$
BERTDS
Group
69.3
65.0
$64.1^{*}$
HaRT
Individual
$\bf{71.1}^{\dagger}$
$\bf{69.8}^{\dagger}$
65.6
$\bf{64.3}^{\dagger}$
GRITage
Ind + Grp
70.8
69.2
64.5
62.7
GRITope
Ind + Grp
70.1
66.5
64.8
61.2
[…]
Table 2 shows the results for stance detection. Both GRIT and HaRT models were fine-tuned with historical language for stance. The PLM trained with individual human context (HaRT) better detects user stance.
Table 2 also compares performances of the models on topic detection (TD) and demographic attribute classification (AC). We use GRIT models pre-trained with 8 training blocks of user texts. Both HaRT and GRIT models have a notion of the user and treat each input as written by a different user. This aspect may aid their performances even for tasks where historical language is unavailable. Even though this information is lacking for TD and AC, fine-tuned HaRT models perform better than the rest. The additional group-attribute pre-training of GRIT models may be introducing noise for document-level tasks, since the results are close to that of fine-tuned HaRT models, yet slightly lower. We can draw parallels between the performance enhancements from GPT-2HLC to HaRT, i.e., a PLM adapted to an out-of-domain corpus (social media) to a PLM trained with individual human context using the same corpus, and between the performance gains from BERTage-MLM to BERTDS, i.e., a PLM adapted to an out-of-domain corpus (blogs) to a PLM trained with group context using the same corpus. Note that Hung et al. (2023) do not report results from BERTage-MLM and BERTDS out-of-domain demographically specialized models. Our results indicate that single-document annotated classification tasks may benefit simply by training PLMs with individual human context.
@@
We also compare the language modeling capability of the various models. Table 3 reports perplexity on the test set of 4438 users from the pre-training data. Multi-task learning is known to impact individual task performance, so we expect a slight dip in perplexity. The results align with our hypothesis and the trend shown in Soni et al. (2022). The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT models perform best while GRIT models result in a slightly lower perplexity than HaRT.
Model Human Context
Test ($ppl$)
GPT-2frozen
None
114.82
GPT-2HLC
None
36.39
HaRT
Individual
28.24
GRITage
Ind + Grp
31.77
GRITope
Ind + Grp
30.32

### v2

# Comparing Pre-trained Human Language Models: Is it Better with Human Context as Groups, Individual Traits, or Both?

@@ Abstract
Incorporating human context into language models is the next frontier for human-centered natural language processing. Currently, two pre-training methods exist: group-wise attributes (e.g., over-45-year-olds) or individual traits. Group attributes are coarse — not all 45-year-olds write the same way — while modeling individual traits allows for a more personalized representation, but requires more complex modeling and data. So far, it is unclear which pre-training approach benefits what tasks. We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on 5 user- and document-level tasks. We find that pre-training with both group and individual features significantly improves the two user-level regression tasks like age estimation and personality assessment. Pre-training on individual users significantly improves the three document-level classification tasks like stance and topic detection. It even does well for downstream tasks without historical user data. Our results suggest both approaches have specific use cases, opening new avenues for human-centered language modeling.
Comparing Pre-trained Human Language Models:
Is it Better with Human Context as Groups, Individual Traits, or Both?
Nikita Soni1, Niranjan Balasubramanian1, H. Andrew Schwartz1,  and Dirk Hovy2 1Stony Brook University, 2Bocconi University {nisoni, niranjan, has}@cs.stonybrook.edu, dirk.hovy@unibocconi.it
@@ 1 Introduction
Language varies between people. To capture this notion, two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language. The first approach focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group. These socio-demographic attributes include age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more. The second approach focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
Figure 1: Pre-training a language model with no human context, with socio-demographic group attribute, with individual traits, and with both group and individual traits.
While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2023): People are not defined by their group membership alone Orlikowski et al. (2023), but individual traits might not generalize. We compare the downstream performance of models pre-trained with the two approaches as well as their combination. We answer the following broader research questions:
(RQ1): How can we incorporate group and/or individual human context into pre-training?
(RQ2): How does pre-training with different human contexts differ in terms of downstream performance for different tasks?
@@
We use the monolingual socio-demographically adapted model from Hung et al. (2023) and the HaRT model from Soni et al. (2022) for group and individual human contexts, respectively. For a model trained with both individual and group human context, we use a multi-task learning PLM based on HaRT to create a model, GRIT, in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness). We test all systems on five user- and document-level tasks.
Note that because we focus on empirically comparing pre-training group and individual traits in PLMs, we cannot compare to large language models like GPT4, which do not support stratification to either attribute via pre-training. Recent studies have explored methodologies like user-adapters (Zhong et al., 2021), and user-centric prompting (Li et al., 2023) to include human context into models. In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts.
@@
We find that PLM pre-trained on individuals and group features enhances 2 user-level regression tasks: age estimation and personality assessment from multiple user documents. These user-level tasks focus on individual people. Our findings, that they are best modeled as a combination of their group attributes and unique characteristics, conforms with the notion that a person is a mix of both. Document-level categorization tasks, like stance detection, are more personal than group based. A PLM pre-trained on individual human context alone improves 3 document-level classification tasks.
@@
By their very nature, models of this kind touch upon sensitive user information. For this reason, we take a responsible release strategy, making only the code for the comparisons publicly available and the exact splits of the TrustPilot and Stance datasets used. We build on top of the publicly available code for HaRT and Hung et al. (2023). We acquired the model and data in a secure manner from the authors of Soni et al. (2022) and TrustPilot data splits from the authors of Hung et al. (2023). For more information about the model and data, see Sections 3 and 4. For a discussion of the ethical implications of the model and data, see Ethical Considerations section.
@@
(1) We compare pre-training several language models with individual users, group socio-demographic features, and both traits.
(2) We evaluate the three specific pre-training strategies on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category prediction).
(3) We find that the two user-level regression tasks perform better when pre-trained with both strategies (GRIT) and the three single document-level tasks perform better on fine-tuning a language model pre-trained with individual context alone (HaRT).
@@ 2 Integrating Human Context in PLMs
For our comparison, we use three systems representing the paradigms of pre-training with human context (Figure 1). We want to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts. As noted earlier, we focus on recent approaches for pre-training language models with the additional human context.
Pre-training with group context.
@@
We build on Hung et al. (2023), a recent model to explore demographic adaptation in transformer-based PLMs. They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author. They use the multilingual reviews dataset with demographic labels from Trustpilot [^1] https://www.trustpilot.com/ Hovy (2015). They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection. We use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4). We compare to the results from Hung et al. (2023) with monolingual BERT on a US-English dataset with out-of-domain demographic (age) specialization. Out-of-domain data is the Blogs authorship corpus Schler et al. (2006), and in-domain means Trustpilot corpus. To be consistent and fair in comparing with other human context pre-training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
Pre-training with individual human context.
@@
Soni et al. (2022) introduced human language modeling (HuLM) in pre-trained LMs. I.e., pre-training a regular language modeling but including a dynamic individual human context vector derived from the authors’ historical texts. This vector captures the human states in which the text was generated. It also adds coherence to texts generated by the same author. They introduce a Human-aware Recurrent Transformer (HaRT), an autoregressive PLM for human language modeling. They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks. We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
Pre-training with both group and individual human context.
We train a PLM to integrate both individual and group human context in language modeling pre-training. We extend HaRT by predicting an additional human group attribute in a dynamic muti-task learning setup similar to Hung et al.. To facilitate comparison, we combine the pre-training inducing the individual human context through the author’s language and inject the group context by predicting a group attribute of the author.
We choose downstream tasks from both prior works (Soni et al., 2022; Hung et al., 2023) to compare our three language models pre-trained with different levels of human contexts: two user-level regression tasks and three single document-level classification task.
@@ 3 Models
3.1 Pre-training with individual human context
@@
where, $w_{t,i}$ is the $i^{th}$ token in the $t^{th}$ block ($B_{t}^{(a)}$) for user $a$. The tokens from the previous blocks are represented using HaRT’s recurrently updated user-state vector. Soni et al. use cross-entropy loss for the HuLM objective.
3.2 Pre-training with group human context
@@
Hung et al. (2023) train for domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, they adopt a dynamic MTL objective for training with group human context. Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions. This approach accounts for the different scales across multiple losses.
@@
3.3 Pre-training with both individual and group human context
@@
GRoup and Individual HaRT is a LM pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup. The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
@@
We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope), respectively. The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$). Similar to Hung et al. (2023), we account for the homoscedastic uncertainty Kendall et al. (2018) of losses. We use the joint loss for a continuous and discrete output as derived in Kendall et al. (2018) and compute our joint objective as follows:
@@ 3.4 Fine-Tuning
3.5 Transfer Learning
We experiment with continuing pre-training GRIT models for each group attribute. To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness). We do the same for GRITope and continued MTL with predicting age.
@@ 4 Experiments
Our research goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined. To this end, we evaluate performances of the models defined in Section 3 on two user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification. We also compare against GPT-2HLC from Soni et al. (2022) as a PLM adapted to the social media domain but devoid of human context.
@@ 4.1 User Level Regression Tasks
We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment Kosinski et al. (2013); Park et al. (2015). These social scientific tasks require predicting continuous outcomes for a user given multiple documents written by them. We use the same data splits as used by Soni et al. (2022) for evaluating HaRT and GPT-2HLC
We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively. We further evaluate these models on the test sets for personality assessment and age estimation, respectively, after continuing training for these tasks, as described in section 3.5.
@@ 4.2 Document-Level Classification Tasks
We use the same subset of the TrustPilot dataset as for topic detection and the same train, development, and test splits from Hung et al. (2023). Given a single review, this task predicts the age group binary label (<35 years old or >45 years old). Age categories are equally represented in each set. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
@@ 5 Results and Discussion
We report results for all the tasks here, discussing their respective impacts from pre-training LMs with individual human context, group context, and both individual and group context.
@@ 5.1 Comparisons Study
User-Level Regression Tasks.
Table 1 shows the two user-level regression task results. We find that just the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality. Additionally, comparing the transfer learning (Section 3.5) results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find the GRIT models to fare better.
Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it is still devoid of human context. HaRT adds pre-training with the individual human context, and GRIT adds pre-training with both group and individual human contexts (Figure 1). As Table 1 shows, there are gains going from GPT-2HLC (no human context) to HaRT (individual human context) and further to GRIT (individual + group human context). This indicates that pre-training PLMs with individual and group human context may benefit multi-document user-level regression tasks like the ones we considered.
@@
GPT-2HLC None 0.839 0.521
HaRT Individual 0.868 0.619
GRITage Ind + Grp 0.890 0.658
GRITope Ind + Grp $0.884$ $0.643$
Table 1: Pearson $r$ for age, disattenuated Pearson $r$ for openness. Pre-training with individual plus group context show benefits in estimating age and assessing personality (openness). Bold = best in column. We find no statistical difference between GRITage and GRITope for the task of age estimation. All other results show statistical significance $p<0.05$ using paired t-test.
Document-Level Classification Tasks.
GPT-2HLC None 68.6 69.8 65.4 63.9
BERTage-MLM Group - 68.4 64.6 $61.9$
BERTDS Group - 69.3 65.0 $64.1$
HaRT Individual $\textbf{71.1}^{*}$ $\textbf{69.8}^{*}$ 65.6 $\textbf{64.3}^{*}$
GRITage Ind + Grp 70.8 69.2 64.5 62.7
GRITope Ind + Grp 70.1 66.5 64.8 61.2
@@
Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute classification (AC). We see that task fine-tuned HaRT (individual human context) models performs better on all tasks..
HaRT models inherently have an additional context of the individual user and do not treat all inputs as if written by the same user. The tasks considered here relate more to personal opinions and preferences, rather than group-level ones. HaRT model is well-suited for incorporating such personalization due to its pre-training with individual human context. Additionally, the results indicate that the models pre-trained with group context (BERTDS) fare well in the group-based tasks of topic detection and age classification. Whereas, the models pre-trained on both individual and group human context (GRIT) appears to bring in noise for the group-based and personal stance detection tasks resulting in a slightly worse performance.
Further, it is important to note that the individual human context (HaRT) derived from the historical tweets from the users in the stance detection dataset provides for a richer human context. Although, the performance is not greatly hurt even if historical language is not available for certain tasks (TD and AC).
@@
We also compare the language modeling capability of the various models. Table 3 reports perplexity on the test set of 4438 users. The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT (individual human context) models perform best while GRIT (individual + group human context) models result in a slightly lower perplexity than HaRT. GRIT models are pre-trained under a multi-task learning setup which most likely is hurting the individual task performances, thus resulting in a slight dip in the perplexity compared to HaRT. Further, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al. (2022).
Model Human Context Test ($ppl$)
GPT-2frozen None 114.82
GPT-2HLC None 36.39
HaRT Individual 28.24
GRITage Ind + Grp 31.77
GRITope Ind + Grp 30.32